# Strict Subgoal Execution: Reliable Long-Horizon Planning in Hierarchical Reinforcement Learning

### v1

@@ Appendix C Experimental Setup
Our proposed framework is designed to be modular and general, enabling integration with a wide range of baseline methods. For comparison, we employ the official codebases provided by the original authors for HIRO, HRAC, HIGL, DHRL, NGTE, and BEAG. All baselines are run using the hyperparameters specified in their respective publications, and conducted on an NVIDIA RTX 3090 GPU with an Intel Xeon Gold 6348 CPU (Ubuntu 20.04). Appendix C.1 provides descriptions of the baseline algorithms along with links to their official code repositories. Appendix C.2 provides specifications of the nine environments shown in Fig.9, including their action and observation spaces, goal configurations, and episode horizons. The SSE-specific hyperparameter configurations for each environment are summarized in AppendixC.3.
@@ C.1 Details of Other Baselines
  • HIRO [40] introduces a hierarchical architecture with relabeling of high-level transitions to account for changing low-level policies, thereby improving off-policy sample efficiency and stability. Open-source code of HIRO is available at https://github.com/watakandai/hiro_pytorch
  • HRAC [16] adds a learned adjacency constraint to ensure subgoal feasibility. It penalizes high-level selections that attempt transitions deemed unreachable within a limited horizon, thereby guiding the agent to learn feasible subgoal structures. Open-source code of HRAC is available at https://github.com/trzhang0116/HRAC
  • HIGL [12] incorporates a coverage-driven and novelty-driven landmark selection strategy. It performs graph-based planning via shortest paths and uses adjacency rewards to guide learning toward under-explored regions. Open-source code of HIGL is available at https://github.com/junsu-kim97/HIGL
  • DHRL [19] constructs a goal graph using Farthest Point Sampling and learns high-level behavior by planning over the graph. It emphasizes temporal abstraction and long-horizon planning via graph traversal and Q-learning. Open-source code of DHRL is available at https://github.com/jayLEE0301/dhrl_official
  • BEAG [15] employs value-function-driven imaginary landmarks to facilitate exploration of unvisited areas. It estimates landmark distances from a learned value function without relying solely on previously visited states, enabling efficient generalization. Open-source code of BEAG is available at https://github.com/ml-postech/BEAG
  • NGTE [20] drives novelty-based exploration by identifying frontier nodes (outposts) and prioritizing expansion toward less-visited regions of the goal graph, encouraging broad and diverse exploration. Open-source code of NGTE is available at https://github.com/ihatebroccoli/NGTE
@@ C.2 Environmental Details
We follow standard benchmarks and configurations widely adopted in prior hierarchical reinforcement learning studies [19, 20, 15], and introduce several new environments designed to evaluate high-level decision-making capabilities such as AntKeyChest or AntDoubleKeyChest. The environments used in our experiments are visualized in Fig. 9, and their characteristics are described in Table 1.
@@
$1$ if key reached
$5$ if goal reached
@@ C.3 Hyperparameter Setup
Table 3 presents environment-specific hyperparameters, including the minimum epsilon threshold $\epsilon_{\text{min}}$, the path refinement scaling factor $c_{\mathrm{dist}}$, the subgoal success threshold $\lambda$, and the grid resolution $d_{\mathcal{G}}$. These values were selected based on targeted parameter searches conducted for each environment.
@@
Scenario $\boldsymbol{\epsilon_{\text{min}}}$ $\boldsymbol{c_{\mathrm{dist}}}$ $\boldsymbol{\lambda}$ $\boldsymbol{d_{\mathcal{G}}}$ $\eta$
@@
U-maze 0.1 5.0 2.0 2 0.1
AntMazeBottleneck 0.1 5.0 2.0 2 0.1
$\pi$-maze 0.1 5.0 2.0 2 0.2
AntMazeComplex 0.1 5.0 2.0 2 0.2
AntMazeDoubleBottleneck 0.1 10.0 2.0 2 0.1
AntKeyChest 0.1 5.0 2.0 2 0.2
AntDoubleKeyChest 0.2 5.0 2.0 2 0.2
@@ D.1 Performance Comparison under Random Goal Setups
In Fig. 10, we compare experimental results under a random goal setting, where the goal is sampled uniformly from the entire valid state space during training, unlike the fixed-goal scenarios discussed in the main text. The reward signal remains sparse, posing a significant challenge for goal discovery and policy optimization. Graph-based methods such as BEAG, NGTE, and SSE, which can autonomously expand their subgoal graph during exploration, maintain strong performance even under random goal sampling, achieving success rates comparable to those in the fixed-goal setting. DHRL also benefits from the randomized goal distribution, particularly in simpler environments like U-maze and AntMazeBottleneck, where the training signal is more frequently encountered. On the other hand, methods such as HIRO, HIGL, and HRAC exhibit limited progress in most environments under random-goal conditions, due to their reliance on fixed-frequency subgoal selection, sparse reward setting, and lack of structured exploration mechanisms. In environments requiring multi-stage reasoning such as AntKeyChest, SSE is the only method that consistently discovers the key and solves the full task. NGTE, while hierarchical and exploration-driven, occasionally learns to acquire the key, but overall exhibits a very low success rate under this random goal configuration.
@@ D.2 Computational Complexity Comparison
Table 4 reports the average per-episode computation time and the number of episodes required to reach a 60% success rate with NGTE and BEAG. We compare against NGTE and BEAG because they are the only baselines that succeed in at least one of the considered fixed-goal settings, making them the most relevant references for evaluating performance and sample efficiency in sparse reward environments. SSE achieves faster convergence across tasks due to its early termination of episodes upon subgoal failure and a reduced number of transitions per episode, which minimizes unnecessary computation and accelerates learning.
This advantage is particularly evident in long-horizon tasks such as AntKeyChest, where conventional HRL methods consume many timesteps even during failed attempts and require extended horizons for high-level planning. In contrast, in shorter-horizon scenarios such as AntMazeBottleneck, the computational benefits of SSE are less significant since subgoal transitions and failure terminations occur less frequently. The operations introduced by SSE, including the construction of $\tilde{s}$, evaluation of subgoal completion, and uniform sampling, are lightweight and have linear time complexity. As a result, SSE imposes minimal computational overhead while maintaining stable training dynamics.
@@ E.1 Trajectory Analysis in Other Environments
Fig. 11 illustrates learning progression in AntMazeBottleneck. At the early stage (a), the agent has not yet discovered feasible subgoals, and the high-level policy $\pi^{h}$ selects subgoals largely at random. During this phase, the exploration policy promotes coverage by gradually expanding into novel regions. In the mid stage (b), the agent fails to traverse the narrow bottleneck, primarily because the low-level policy $\pi^{l}$ lacks sufficient experience in that region and the associated failure statistics $\mathrm{ratio}_{\mathrm{fail}}$ remain underrepresented. By the final stage (c), SSE successfully leverages failure-aware refinement and subgoal selection, allowing $\pi^{h}$ to navigate through the bottleneck and reach the goal reliably.
@@
A similar trend is observed in AntMazeDoubleBottleneck as Fig. 12. Initially (a), most regions are unexplored, and subgoal selection remains uninformed. In the intermediate phase (b), the agent again struggles with the second bottleneck due to insufficient failure feedback. As training progresses (c), subgoals selected by $\pi^{h}$ become more effective, and the updated path refinement guides the agent through safer routes, enabling successful traversal of both bottlenecks and completion of the long-horizon task.
@@ E.2 Additional Ablation Studies
Fig. 14 presents the performance of SSE variants with key components ablated. Specifically, we evaluate the following variants: SSE without path refinement (PR), without $\tilde{g}_{\mathrm{novel}}$, $\tilde{g}_{\max}$, or goal $g$ in $\pi^{\mathrm{exp}}$, without the entire $\pi^{\mathrm{exp}}$, and with a fixed low-level step size (FLS).
In complex tasks like AntKeyChest and AntDoubleKeyChest, removing max-value guidance ($\tilde{g}_{\max}$) or novelty-based exploration ($\tilde{g}_{\mathrm{novel}}$) leads to notable performance drops, due to poor goal-space coverage and delayed reward discovery. In contrast, simpler tasks like AntMazeBottleneck are less sensitive to these components. Path refinement similarly plays a larger role in long-horizon settings by mitigating unreliable transitions, while offering little benefit in short tasks with fewer failures. Removing the explicit goal signal (SSE w/o $g$) or exploration policy (SSE w/o $\pi^{\mathrm{exp}}$) consistently harms performance across all environments, confirming the importance of goal-conditioned reasoning and structured exploration. Lastly, replacing SSE’s flexible execution with a fixed-length scheduler (SSE w FLS) shows little effect in short tasks, but severely limits learning in long-horizon, sparse-reward environments like AntDoubleKeyChest.
@@
Fig. 15 illustrates the effect of the path refinement scaling factor $c_{\mathrm{dist}}$ on performance across different environments. In short-horizon tasks such as AntMazeBottleneck, the sensitivity to $c_{\mathrm{dist}}$ is relatively low, and changes in its value have little impact on overall performance.
In contrast, for longer-horizon tasks such as AntKeyChest and AntDoubleKeyChest, the effect of $c_{\mathrm{dist}}$ becomes more significant. When the value increases from 1 to 5, the agent benefits from more reliable path selection due to moderate penalty scaling, which helps avoid failure-prone regions while maintaining reachability. However, setting $c_{\mathrm{dist}}$ to overly large values such as 10 or 20 leads to frequent detours and unstable planning, ultimately degrading both learning efficiency and task success.
@@
Fig. 16 illustrates how the grid resolution $d_{\mathcal{G}}$, used to partition the goal space for novelty identification, affects exploration. A coarse grid with $d_{\mathcal{G}}=4$ fails to capture fine-grained novelty, limiting the ability to distinguish under-visited regions. This shortcoming is pronounced in environments with narrow passages, such as AntMazeBottleneck, where structural bottlenecks are poorly reflected. In contrast, $d_{\mathcal{G}}=1$ creates overly fine partitions, fragmenting the goal space and introducing noise into failure statistics, which leads to unstable path updates and slower learning.
These results highlight the need for an appropriate balance between granularity and stability. Empirically, we use $d_{\mathcal{G}}=2$ for AntMaze environments to ensure reliable exploration while avoiding excessive updates. For Reacher environments, which feature a smaller and more structured goal space, a finer resolution of $d_{\mathcal{G}}=1$ is applied to retain sufficient detail without incurring instability.

### v2

@@ Appendix C Experimental Setup
Our proposed framework is designed to be modular and general, enabling integration with a wide range of baseline methods. For comparison, we employ the official codebases provided by the original authors for HIRO, HRAC, HIGL, DHRL, NGTE, and BEAG. All baselines are run using the hyperparameters specified in their respective publications, and conducted on an NVIDIA RTX 3090 GPU with an Intel Xeon Gold 6348 CPU (Ubuntu 20.04). Appendix C.1 provides descriptions of the baseline algorithms along with links to their official code repositories. Appendix C.2 provides the specifications for the nine environments shown in Fig.9, including their action and observation spaces, goal configurations, and episode horizons. The SSE-specific hyperparameter configurations for each environment are summarized in AppendixC.3.
@@ C.1 Details of Other Baselines
  • HIRO (Nachum et al., 2018b) introduces a hierarchical architecture with relabeling of high-level transitions to account for changing low-level policies, thereby improving off-policy sample efficiency and stability. Open-source code of HIRO is available at https://github.com/watakandai/hiro_pytorch
  • HRAC (Zhang et al., 2020) adds a learned adjacency constraint to ensure subgoal feasibility. It penalizes high-level selections that attempt transitions deemed unreachable within a limited horizon, thereby guiding the agent to learn feasible subgoal structures. Open-source code of HRAC is available at https://github.com/trzhang0116/HRAC
  • HIGL (Kim et al., 2021) incorporates a coverage-driven and novelty-driven landmark selection strategy. It performs graph-based planning via shortest paths and uses adjacency rewards to guide learning toward under-explored regions. Open-source code of HIGL is available at https://github.com/junsu-kim97/HIGL
  • DHRL (Lee et al., 2022) constructs a goal graph using Farthest Point Sampling and learns high-level behavior by planning over the graph. It emphasizes temporal abstraction and long-horizon planning via graph traversal and Q-learning. Open-source code of DHRL is available at https://github.com/jayLEE0301/dhrl_official
  • BEAG (Yoon et al., 2024) employs value-function-driven imaginary landmarks to facilitate exploration of unvisited areas. It estimates landmark distances from a learned value function without relying solely on previously visited states, enabling efficient generalization. Open-source code of BEAG is available at https://github.com/ml-postech/BEAG
  • NGTE (Park et al., 2024) drives novelty-based exploration by identifying frontier nodes (outposts) and prioritizing expansion toward less-visited regions of the goal graph, encouraging broad and diverse exploration. Open-source code of NGTE is available at https://github.com/ihatebroccoli/NGTE
@@ C.2 Environmental Details
We follow standard benchmarks and configurations widely adopted in prior hierarchical reinforcement learning studies (Lee et al., 2022; Park et al., 2024; Yoon et al., 2024), and introduce several new environments designed to evaluate high-level decision-making capabilities such as AntKeyChest or AntDoubleKeyChest. The environments used in our experiments are visualized in Fig. 9, and their characteristics are described in Table 1.
@@
11 if key reached
55 if goal reached
@@ C.3 Hyperparameter Setup
Table 3 presents environment-specific hyperparameters, including the minimum epsilon threshold $\epsilon_{\text{min}}$, the path refinement scaling factor $c_{\mathrm{dist}}$, the reachability check threshold $\lambda$, and the grid resolution $d_{\mathcal{G}}$. As mentioned in Section 2.2, the threshold $\lambda$ is the standard, environment-provided value used to check for low-level reachability (see equation 8), which is also used consistently across baselines. These values were selected based on targeted parameter searches conducted for each environment.
@@
Scenario $\bm{\epsilon_{\text{min}}}$ $\bm{c_{\mathrm{dist}}}$ $\bm{\lambda}$ $\bm{d_{\mathcal{G}}}$ $\eta$
@@
U-maze 0.1 5.0 0.5 2 0.1
AntMazeBottleneck 0.1 5.0 0.5 2 0.1
$\pi$-maze 0.1 5.0 0.5 2 0.2
AntMazeComplex 0.1 5.0 0.5 2 0.2
AntMazeDoubleBottleneck 0.1 10.0 0.5 2 0.1
AntKeyChest 0.1 5.0 0.5 2 0.2
AntDoubleKeyChest 0.2 5.0 0.5 2 0.2
@@ D.1 Performance Comparison under Random Goal Setups
Figure 10 presents experimental results under a random goal setting where goals are sampled uniformly from the entire valid state space during training. This setup contrasts with the fixed-goal scenarios discussed in the main text. The reward signal remains sparse and poses a significant challenge for goal discovery and policy optimization. Graph-based methods such as BEAG, NGTE, and SSE maintain strong performance even under random goal sampling. These methods autonomously expand their subgoal graph during exploration and achieve success rates comparable to those in the fixed-goal setting. DHRL also benefits from the randomized goal distribution, particularly in simpler environments like U-maze and AntMazeBottleneck where the agent encounters training signals more frequently. Conversely, methods such as HIRO, HIGL, and HRAC exhibit limited progress in most environments under random-goal conditions. This performance drop stems from their reliance on fixed-frequency subgoal selection and the lack of structured exploration mechanisms suitable for sparse reward settings. In environments requiring multi-stage reasoning such as AntKeyChest, SSE is the only method that consistently discovers the key and solves the full task. While NGTE is hierarchical and exploration-driven, it only occasionally learns to acquire the key and exhibits a very low overall success rate under this configuration.
@@ D.2 Computational Complexity Comparison
Table 4 reports the average per-episode computation time and the number of episodes required to reach a 60% success rate. We compare SSE against NGTE and BEAG as they are the only baselines that succeed in at least one of the considered fixed-goal settings. This makes them the most relevant references for evaluating performance and sample efficiency in sparse reward environments. SSE achieves faster convergence across tasks due to its early termination of episodes upon subgoal failure and a reduced number of transitions per episode. These features minimize unnecessary computation and accelerate learning.
This advantage is particularly evident in long-horizon tasks such as AntKeyChest. In these tasks, conventional HRL methods consume many timesteps even during failed attempts and require extended horizons for high-level planning. In shorter-horizon scenarios like AntMazeBottleneck, the computational benefits of SSE are less significant since subgoal transitions and failure terminations occur less frequently. The operations introduced by SSE, including the evaluation of subgoal completion and uniform sampling, are lightweight and have linear time complexity. Consequently, SSE imposes minimal computational overhead while maintaining stable training dynamics.
@@ E.1 Trajectory Analysis in Other Environments
Fig. 11 illustrates learning progression in AntMazeBottleneck. In the early stage (a), the agent has not yet discovered feasible subgoals, causing the high-level policy $\pi^{h}$ to select subgoals largely at random. During this phase, the exploration policy promotes coverage by gradually expanding into novel regions. In the mid stage (b), the agent fails to traverse the narrow bottleneck, primarily because the low-level policy $\pi^{l}$ lacks sufficient experience in that region and the associated failure statistics $\rho_{\mathrm{f}}$ remain underrepresented. By the final stage (c), SSE successfully leverages failure-aware refinement and subgoal selection. This allows $\pi^{h}$ to navigate through the bottleneck and reach the goal reliably.
@@
A similar trend is observed in AntMazeDoubleBottleneck, as shown in Fig. 12. Initially (a), most regions are unexplored, and subgoal selection remains uninformed. In the intermediate phase (b), the agent again struggles with the second bottleneck due to insufficient failure feedback. As training progresses (c), subgoals selected by $\pi^{h}$ become more effective. The updated path refinement guides the agent through safer routes, enabling successful traversal of both bottlenecks and completion of the long-horizon task.
@@ E.2 Additional Ablation Studies
Fig. 14 presents the performance of SSE variants with key components ablated. We evaluate four variants: (1) SSE w/o $\mathcal{B}_{F}^{h}$ which replaces FER with a standard replay buffer, (2) SSE with HER which uses a conventional fixed-step policy, (3) SSE w/o $\pi^{\mathrm{exp}}$ which disables decoupled exploration, and (4) SSE w/o PR which omits path refinement. Removing path refinement degrades performance in long-horizon settings where mitigating unreliable transitions is crucial. Similarly, disabling the exploration policy consistently impairs goal-space coverage across all environments. Critically, SSE with HER and SSE w/o $\mathcal{B}_{F}^{h}$ fail in complex tasks like AntDoubleKeyChest despite achieving minor success in simpler maps. This failure demonstrates that conventional HER offers insufficient feasibility signals. Furthermore, it confirms that FER is vital for providing stop-on-failure feedback to teach the high-level policy to avoid unreliable subgoals. These components collectively ensure the decisive signals required for long-horizon planning.
@@
Fig. 15 shows how the scaling factor $c_{\mathrm{dist}}$ affects performance. This parameter controls the trade-off between path efficiency and safety. In simpler tasks like AntMazeBottleneck, its impact is minimal. In more complex, long-horizon tasks, its role becomes more pronounced. A moderate value, such as $c_{\mathrm{dist}}=5$, provides an effective balance, guiding the planner away from failure-prone regions without being overly conservative. In contrast, excessively high values (e.g., 10 or 20) can lead to inefficient detours, while a value of 1 may not sufficiently penalize risky paths. Overall, the performance remains high for values between 2 and 10, indicating a wide effective range.
@@
Fig. 16 illustrates the effect of grid resolution $d_{\mathcal{G}}$. This parameter balances the granularity of novelty detection with the stability of failure statistics. The analysis shows that performance is not highly sensitive to this parameter within the tested range. For instance, in AntMaze environments, there is little significant difference in performance for $d_{\mathcal{G}}$ values of 1, 2, and 4. A coarse grid ($d_{\mathcal{G}}=4$) may merge distinct regions, while an overly fine grid ($d_{\mathcal{G}}=1$) can make failure statistics less reliable. Thus, we find that a moderate resolution ($d_{\mathcal{G}}=2$ for AntMaze) provides a suitable balance. The optimal choice is dependent on the environment’s scale, with smaller, structured spaces like Reacher benefiting from a finer grid ($d_{\mathcal{G}}=1$).
@@
Fig. 17 shows how performance varies with the exploration ratio $\eta$, which controls the classic exploration-exploitation trade-off. In simple environments like AntMazeBottleneck, performance is consistent across a wide range of $\eta$ values. In long-horizon tasks like AntDoubleKeyChest that require discovering intermediate objectives, the influence of $\eta$ is naturally greater, but effective performance can be achieved with straightforward tuning. As shown in our experiments, setting the exploration ratio within a small range of 0.1 to 0.2 is sufficient to achieve strong performance across all tasks. This indicates that while the balance is important, extensive tuning of this hyperparameter is not required.